# Minority-Focused Text-to-Image Generation via Prompt Optimization

### v2

@@ Abstract
We investigate the generation of minority samples using pretrained text-to-image (T2I) latent diffusion models. Minority instances, in the context of T2I generation, can be defined as ones living on low-density regions of text-conditional data distributions. They are valuable for various applications of modern T2I generators, such as data augmentation and creative AI. Unfortunately, existing pretrained T2I diffusion models primarily focus on high-density regions, largely due to the influence of guided samplers (like CFG) that are essential for producing high-quality generations. To address this, we present a novel framework to counter the high-density-focus of T2I diffusion models. Specifically, we first develop an online prompt optimization framework that can encourage the emergence of desired properties during inference while preserving semantic contents of user-provided prompts. We subsequently tailor this generic prompt optimizer into a specialized solver that promotes the generation of minority features by incorporating a carefully-crafted likelihood objective. Our comprehensive experiments, conducted across various types of T2I models, demonstrate that our approach significantly enhances the capability to produce high-quality minority instances compared to existing samplers.
@@ 1 Introduction
In this work, we present a novel approach dubbed as MinorityPrompt that counteracts the high-density bias of T2I samplers to improve their capability of minority generation. Our framework is built upon the concept of prompt optimization, an intuitive technique that exhibits strong performance in enhancing T2I diffusion models for various tasks [14, 6, 36]. Unlike existing T2I-based online prompt-tuning methods that modify the entire input prompts (e.g., by optimizing their text-embeddings during inference), our approach updates the prompts in a selective fashion to preserve the intended semantics while encouraging generations of unique low-density features. Specifically during inference, we incorporate learnable tokens into the input prompts, e.g., by appending them to the end of the text. We then adjust the embeddings of these tokens across sampling timesteps, targetting the minimization of a likelihood metric designed to capture the uniqueness of noisy intermediate samples. See Fig. 2 for an overview. An additional benefit of our token-based approach is that it offers enhanced semantic controllability, enabling users to express specific desired semantics in generated samples by selecting appropriate initialization words for the learnable token embeddings. To further improve the performance of our sampler, we provide new design choices that can be synergistically employed with our approach for T2I minority generation. Comprehensive experiments validate that our method can significantly improve the ability of creating minority instances of modern widely-adopted T2I models (including Stable Diffusion (SD) [41]) with minimal compromise in sample quality and text-image alignment. In addition, we emphasize that our framework can work on distilled backbones like SDXL-Lightning [29], which demonstrates its robustness and practical relevance. As an additional application, we explore the potential of our prompt optimization framework to improve the diversity of T2I models, further exhibiting its versatility as a general-purpose solver applicable across various tasks.
Given that our prompt optimization is performed in an online manner, does not require expensive fine-tuning of T2I models, and is entirely self-contained, i.e., implementable solely with a pretrained T2I model, we believe our approach open a new avenue for creative AI, emphasizing the practical relevance of our framework.
@@ 2 Related Work
The generation of minority samples has been explored in a range of different scenarios and generative frameworks [56, 31, 44, 38, 22, 51, 52]. However, significant progress has been recently made with the introduction of diffusion models, due to their ability to faithfully capture data distributions [44, 51, 52]. As an initial effort, [44] incorporate separately-trained classifiers into the sampling process of diffusion models to yield guidance for low-density regions. The approach by [51] shares similar intuition of integrating an additional classifier into the reverse process for low-density guidance. A limitation is that their methods rely upon external classifiers that are often difficult to obtain, especially for large-scale datasets such as T2I benchmarks [43]. The challenge was recently addressed by [52] where the authors develop a self-contained minority sampler that works without expensive extra components (such as classifiers). However, their method is tailored for canonical image benchmarks (like LSUN [55] and ImageNet [9]) and exhibits limited performance gain in more challenging scenarios like T2I generation. Moreover, none of these approaches have explored the dimension of prompt optimization specifically for minority generation, which is the central focus of our framework.
@@ 3.1 Background and preliminaries
$\max_{\boldsymbol{\theta}}\mathbb{E}_{{{\boldsymbol{z}}_{0},y,{\boldsymbol{\epsilon}}\sim{\cal N}({\boldsymbol{0}},{\boldsymbol{I}}),t\sim\text{Unif}\{1,\dots,T\}}}[\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\boldsymbol{\theta}}({\boldsymbol{z}}_{t},{\cal C})\|_{2}^{2}],$
@@ 4.2 Results
where ${\cal B}$ is the batch size, and $\{\boldsymbol{z}_{t}^{(i)}\}_{i=1}^{\cal B}$ denotes the noisy intermediate samples in the batch given at time $t$. We found that incorporating this objective into Eq. (4) yields impressive results, even rivaling the state-of-the-art diverse sampler [42]; see Tab. 3 for details.
@@ 7.2 Implementation details
Other details. Our implementation is based on PyTorch [37], and experiments were performed on twin NVIDIA A100 GPUs. Code is available at https://github.com/anonymous5293/MinorityPrompt.
@@ 9.3 Additional generated samples
To facilitate a more comprehensive qualitative comparison among the samplers, we provide an extensive showcase of generated samples for all the focused T2I pretrained models. See Figures 79 for details.

### v3

@@ Abstract
We investigate the generation of minority samples using pretrained text-to-image (T2I) latent diffusion models. Minority instances, in the context of T2I generation, can be defined as ones living on low-density regions of text-conditional data distributions. They are valuable for various applications of modern T2I generators, such as data augmentation and creative AI. Unfortunately, existing pretrained T2I diffusion models primarily focus on high-density regions, largely due to the influence of guided samplers (like CFG) that are essential for high-quality generation. To address this, we present a novel framework to counter the high-density-focus of T2I diffusion models. Specifically, we first develop an online prompt optimization framework that encourages emergence of desired properties during inference while preserving semantic contents of user-provided prompts. We subsequently tailor this generic prompt optimizer into a specialized solver that promotes generation of minority features by incorporating a carefully-crafted likelihood objective. Extensive experiments conducted across various types of T2I models demonstrate that our approach significantly enhances the capability to produce high-quality minority instances compared to existing samplers. Code is available at https://github.com/soobin-um/MinorityPrompt.
@@ 1 Introduction
In this work, we present a novel approach dubbed as MinorityPrompt that counteracts the high-density bias of T2I samplers to improve their capability of minority generation. Our framework is built upon the concept of prompt optimization, an intuitive technique that exhibits strong performance in enhancing T2I diffusion models for various tasks [14, 6, 36]. Unlike existing T2I-based online prompt-tuning methods that modify the entire input prompts (e.g., by optimizing their text-embeddings during inference), our approach updates the prompts in a selective fashion to preserve the intended semantics while encouraging generations of unique low-density features.
Specifically during inference, we incorporate learnable tokens into the input prompts, e.g., by appending them to the end of the text. The embeddings of these tokens are iteratively refined across sampling timesteps to optimize the proposed objective for minority generation, which approximates the likelihood of noisy intermediate samples in T2I generation. See Fig. 2 for an overview. We highlight that our prompt optimization framework is versatile, i.e., it can be applied to various tasks with distinct optimization objectives beyond minority generation.
Comprehensive experiments validate that our method can significantly improve the ability of creating minority instances of modern widely-adopted T2I models (including Stable Diffusion (SD) [41]) with minimal compromise in sample quality and text-image alignment. In addition, we emphasize that our framework can work on distilled backbones like SDXL-Lightning [29], which demonstrates its robustness and practical relevance. As an additional application, we explore the potential of our prompt optimization framework to improve the diversity of T2I models, further exhibiting its versatility as a general-purpose solver applicable across various tasks.
Our key contributions are summarized as follows:
  • We propose a token-based online prompt optimization framework that iteratively updates learnable tokens during inference, achieving superior text-image alignment over existing online prompt tuners.
  • We develop a novel objective for minority sampling in the T2I context, which mathematically approximates the target log-likelihood in T2I generation.
  • We empirically demonstrate that our approach achieves state-of-the-art performance in generating minority samples for T2I generation.
@@ 2 Related Work
The generation of minority samples has been explored in a range of different scenarios and generative frameworks [56, 31, 44, 38, 22, 51, 52]. However, significant progress has been recently made with the introduction of diffusion models, due to their ability to faithfully capture data distributions [44, 51, 52]. As an initial effort, [44] incorporate separately-trained classifiers into the sampling process of diffusion models to yield guidance for low-density regions. The approach by [51] shares similar intuition of integrating an additional classifier into the reverse process for low-density guidance. A limitation is that their methods rely upon external classifiers that are often difficult to obtain, especially for large-scale datasets such as T2I benchmarks [43]. The challenge was recently addressed by [52] where the authors develop a self-contained minority sampler that works without expensive extra components (such as classifiers). However, their method is tailored for canonical image benchmarks (like LSUN [55] and ImageNet [9]) and exhibits limited performance gain in more challenging scenarios like T2I generation.
@@ 3.1 Background and preliminaries
$\min_{\boldsymbol{\theta}}\mathbb{E}_{{{\boldsymbol{z}}_{0},y,{\boldsymbol{\epsilon}}\sim{\cal N}({\boldsymbol{0}},{\boldsymbol{I}}),t\sim\text{Unif}\{1,\dots,T\}}}[\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\boldsymbol{\theta}}({\boldsymbol{z}}_{t},{\cal C})\|_{2}^{2}],$
@@ 4.2 Results
where ${\cal B}$ is the batch size, and $\{\boldsymbol{z}_{t}^{(i)}\}_{i=1}^{\cal B}$ denotes the intermediate samples in the batch at $t$. We found that incorporating Eq. (9) into Eq. (4) yields impressive results, even rivaling the state-of-the-art diverse sampler [42] (see Tab. 3). We leave generated samples in the supplementary.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) under Grants RS-2024-00336454 and RS-2023-00262527, and by Basic Science Research Program through the NRF funded by the Ministry of Education (RS-2024-00406726).
@@ 7.2 Implementation details
Other details. Our implementation is based on PyTorch [37], and experiments were performed on twin NVIDIA A100 GPUs. Code is available at https://github.com/soobin-um/MinorityPrompt.
@@ 9.2 User preference study
Figure 10: Generated samples from diversity-enhancing approaches on SDv1.5. Samples generated using three distinct methods: (i) DDIM [46]; (ii) CADS [42]; (iii) Ours (i.e., Eq. (9)). All samplers shared the same initial noise for generation.
@@ 9.3 Additional generated samples
To facilitate a more comprehensive qualitative comparison among the samplers, we provide an extensive showcase of generated samples for all the focused T2I pretrained models. See Figures 7-9 for details. In addition, we exhibit samples generated using various diversity-focused approaches (discussed at the end of Sec. 4.2); see Fig. 10 for further details.